# Regularized Estimation of Loading Matrix in Factor Models for High-Dimensional Time Series

### v2

@@ Abstract
High-dimensional data analysis using traditional models suffers from overparameterization. Two types of techniques are commonly used to reduce the number of parameters – regularization and dimension reduction. In this project, we combine them by imposing a sparse factor structure and propose a regularized estimator to further reduce the number of parameters in factor models. A challenge limiting the widespread application of factor models is that factors are hard to interpret, as both factors and the loading matrix are unobserved. To address this, we introduce a penalty term when estimating the loading matrix for a sparse estimate. As a result, each factor only drives a smaller subset of time series that exhibit the strongest correlation, improving the factor interpretability. The theoretical properties of the proposed estimator are investigated. The simulation results are presented to confirm that our algorithm performs well. We apply our method to Hawaii tourism data.
Keywords: Regularization; dimension reduction; ADMM algorithm.
@@ 1 Introduction
High-dimensional time series data are widely observed in various disciplines, including finance, economics, business, and medical science. However, when analyzing high-dimensional data, traditional models, such as vector autoregressive models and regression models, suffer from overparameterization as noted in Lütkepohl, (2005) and Chang et al., (2015). Consequently, traditional methods designed for these models are often not satisfactory in terms of prediction accuracy and model interpretability (Tibshirani,, 1996). To address this problem, two types of methods are commonly used to reduce the number of parameters — regularization and dimension reduction (Liu and Zhang,, 2022). Regularization methods resort to a sparsity assumption on the coefficient matrices and incorporate a penalty in the optimization function to obtain meaningful estimators; see examples Basu and Michailidis, (2015) and Medeiros and Mendes, (2016), while dimension reduction assumes that the high-dimensional data can be represented by a low-dimensional process. Among various dimension reduction tools, factor analysis is a popular choice and has been extensively investigated in the literature (Peña and Box,, 1987; Li and Li,, 1996; Bai and Ng,, 2002; Stock and Watson, 2002a, ; Diebold and Li,, 2006; Lam et al.,, 2011; Ando and Bai,, 2017). In this paper, we combine both methods by imposing a sparse factor structure on the observed data and propose a regularized estimator for the loading matrix to further reduce the number of parameters.
@@
Factor analysis has been an active research area in statistics and econometrics with a comprehensive theoretical framework (Bai and Ng,, 2002; Bai,, 2003; Lam et al.,, 2011; Lam and Yao,, 2012) and numerous extensions, including factor models with structural breaks (Chen,, 2015; Baltagi et al.,, 2017; Barigozzi et al.,, 2018; Baltagi et al.,, 2021), threshold mechanism (Massacci,, 2017; Liu and Chen,, 2020), regime switching (Liu and Chen,, 2016), and constraints (Chen et al.,, 2020), and factor models for matrix data (Wang et al.,, 2019) and tensor data (Chen et al.,, 2022; Han et al.,, 2022; Han et al., 2024a, ; Han et al., 2024b, ). One issue of factor analysis is that factors are difficult to interpret as both factors and the loading matrix are latent. Ideally, observed variables can be used to proxy unobserved common factors. Bai and Ng, 2006b proposed statistical tests to evaluate the latent and observed factors. However, identifying suitable proxy variables is often challenging, especially when dealing with new data or data that are not well understood. Another solution is to interpret factors via loadings. In this paper, we incorporate a penalty function into the estimation of the loading matrix to shrink certain loadings to zero. This encourages each factor to load on a smaller subset of time series that exhibit the strongest correlation and thus enhances interpretability, thereby providing guidance on how to find meaningful proxy variables.
@@
Two widely used statistical methods aimed at improving the interpretability of factor models are particularly worth mentioning — varimax rotation and sparse principal component analysis (SPCA). It is well known that the loading matrix is not uniquely defined and can rotate in the loading space, which is obtained by principal component analysis (PCA) or other matrix decomposition methods. Varimax seeks the rotation that maximizes the variance of the squared elements in the loading matrix. Hence, loadings yielded by varimax are either very large in magnitude or close to 0. Compared to varimax, the method we propose has two key advantages. First, while varimax produces estimated loadings that are near-zero, but our method ensures an estimate with zero elements, enhancing factor interpretability. Second, our method searches for the solution in a more general setting, yielding a more sparse loading matrix estimate. Unlike varimax, which keeps the loading space invariant, we allow the estimator to deviate slightly from the loading space obtained from PCA or its variants. SPCA is an extension of PCA (Zou et al.,, 2006) and requires the principal components to be orthogonal. In contrast to SPCA, our method relaxes this assumption, leading to a more sparse loading matrix and improving the interpretability of the factor model.
@@
Regularization approaches are well developed and have been applied to various models, such as regression models (Fan and Lv,, 2010; Huang et al.,, 2012), graph models (Meinshausen and Bühlmann,, 2006; Liu et al.,, 2009) and model-based clustering for regression coefficients (Ma and Huang,, 2017; Wang et al.,, 2023). These approaches employ different penalty functions to impose the sparsity of the parameter space. LASSO, introduced by Tibshirani, (1996), using an $L_{1}$ penalized likelihood for linear regression models, drew extensive attention, leading to numerous advances in its applications. Zou, (2006) proposed adaptive LASSO to improve the performance of LASSO. Subsequently, other sparsity-inducing penalties, such as smoothly clipped absolute deviation (SCAD) (Fan and Li,, 2001) and minimax concave penalty (MCP) (Zhang,, 2010), were also considered. Estimators obtained by SCAD and MCP both enjoy the oracle properties, meaning they perform as well as if the true underlying model were known in advance (Zou,, 2006), which is theoretically appealing (Bertsimas et al.,, 2020). However, regularization methods have not been extensively studied in the context of factor models. Uematsu and Yamagata, 2022a and Uematsu and Yamagata, 2022b explored the sparsity-induced weak factor models. As mentioned earlier, the loading matrix can rotate in the loading space. Uematsu and Yamagata, 2022a first selected a specific rotation, which makes the factor covariance matrix an identity matrix and columns of the loading matrix orthogonal, and then imposed the sparsity assumption on that particular loading matrix. However, the restrictions that factors are independent and the loading matrix is orthogonal are often too strong for empirical analysis. For example, the Fama-French three-factor models include an overall market factor and factors related to firm size (SMB) and book-to-market equity (HML). Empirical evidence shows that the correlation between the overall market factor and HML is significantly different from 0, which is -0.38 in Fama and French, (1993), -0.622 in Durand et al., (2011), -0.30 in Fama and French, (2015). Another example is the geographic location (i.e. Europe, Asia, Africa, etc) and the economic condition including the level of economic development (i.e. advanced vs emerging), which are considered as two important determinants for international business activities (Kose et al.,, 2003; Sanyal,, 2005; Fernhaber et al.,, 2008). Since Europe has the most advanced economies and Africa has the most emerging economies, it is not appropriate to assume that the loadings of these two factors are orthogonal. THese restrictions are often too strong for empirical analysis. In this paper, we relax these restrictions and explore the entire loading space to identify the most sparse loading matrix, which is more general and leads to a more sparse estimate. Numerical experiments and real data analysis in Section 5 and Section 6 show that our method outperforms and works very well.
The primary contribution of this paper is to propose a regularized estimator for factor models in a general setting. Unlike the existing approaches, our method does not assume an orthogonal loading matrix and independent factors. Consequently, our proposed method may improve the interpretation of factor models. Although factor analysis for time series and regularization methods are two well-studied research areas, their intersection— regularized estimation for factor models remains largely unexplored. In this paper, we develop an algorithm based on the alternating direction method of multipliers (ADMM) algorithm (Boyd et al.,, 2011) with the MCP to obtain a sparse loading matrix. Both numerical simulations and real data examples demonstrate the effectiveness of our method in producing a sparse loading matrix estimate. The analysis on Hawaii tourism data illustrates that our regularized estimators provide insightful interpretations of the latent factors. Furthermore, we establish the consistency of the proposed estimators and study the convergence rate of the estimators, which depends on the sample size, dimension, factor strength, and the sparsity level of the loading matrix. Under mild conditions, the oracle properties remain valid.
@@ 2 Model
We introduce some notations first. For a vector ${\mathbf{z}}$, we use $z_{i}$ to denote its $i$-th element. For a $p_{1}\times p_{2}$ matrix ${\mathbf{Z}}$, its ($i,j$)-th element is denoted by $z_{ij}$ and its $i$-th column is denoted by ${\mathbf{z}}_{i}$. Furthermore, we use ${\cal M}({\mathbf{Z}})$ to denote the space spanned by the columns of ${\mathbf{Z}}$. Let $\|{\mathbf{Z}}\|_{F}$ be the Frobenius norm of ${\mathbf{Z}}$, where $\|{\mathbf{Z}}\|_{F}=\sqrt{\sum_{i=1}^{p_{1}}\sum_{j=1}^{p_{2}}z_{ij}^{2}}$, $\|{\mathbf{Z}}\|_{2}$ be the $L$-2 norm of ${\mathbf{Z}}$, where $\|{\mathbf{Z}}\|_{2}=\left[\lambda_{\max}({\mathbf{Z}}^{\top}{\mathbf{Z}})\right]^{1/2}$ and $\lambda_{\max}(\cdot)$ is the maximum eigenvalue of a square matrix, and $\|{\mathbf{Z}}\|_{\min}$ is the nonzero minimum singular value of ${\mathbf{Z}}$. We also define $L_{1}$, $L_{\infty}$, and max norm of a matrix as follows: $\|{\bf Z}\|_{1}=\max_{1\leq j\leq p_{2}}\sum_{i=1}^{p_{1}}|z_{ij}|$, $\|{\bf Z}\|_{\infty}=\max_{1\leq i\leq p_{1}}\sum_{j=1}^{p_{2}}|z_{ij}|$, and $\|{\bf Z}\|_{\max}=\max_{ij}|z_{ij}|$. We follow Cape et al., (2019) and define two-to-infinity norm as $\|{\bf z}\|_{\infty}=\max_{i}|z_{i}|$ for a vector ${\mathbf{z}}$ and as $\|{\bf Z}\|_{2\rightarrow\infty}=\sup_{\|{\mathbf{x}}\|_{2}=1}\|{\bf Z}{\bf x}\|_{\infty}$ for a matrix ${\mathbf{Z}}$. For a scalar $z$, $|z|$ denotes its absolute value; for a set $\mathcal{Z}$, $|\mathcal{Z}|$ is the number of elements in the set; for a vector ${\mathbf{z}}$ or a matrix ${\mathbf{Z}}$, $|{\mathbf{z}}|$ or $|{\mathbf{Z}}|$ is the number of its nonzero elements. For any $\{a_{n}\}$ and $\{b_{n}\}$, “$a_{n}\asymp b_{n}$” means $\lim_{n\to\infty}a_{n}/b_{n}=c$, where $c$ is a positive constant, and “$a_{n}\gtrsim b_{n}$” means $a_{n}^{-1}b_{n}=o(1)$.
@@
${\mathbf{x}}_{t}={\mathbf{A}}{\mathbf{f}}_{t}+\mbox{\boldmath$\varepsilon$}_{t},$ (1)
[…]
@@
${\mathbf{x}}_{t}={\mathbf{A}}^{s}{\mathbf{f}}_{t}^{s}+\mbox{\boldmath$\varepsilon$}_{t},$ (2)
@@
${\mathbf{x}}_{t}={\mathbf{Q}}{\mathbf{z}}_{t}+\mbox{\boldmath$\varepsilon$}_{t},$ (3)
where ${\mathbf{q}}_{1}=\frac{{\mathbf{a}}_{1}^{s}}{\|{\mathbf{a}}_{1}^{s}\|_{2}}$ and ${\mathbf{q}}_{i}=\frac{{\mathbf{a}}_{i}^{s}}{\|{\mathbf{P}}_{i}{\mathbf{a}}_{i}^{s}\|_{2}}$, for $i=2,\ldots,r$, where ${\mathbf{P}}_{i}={\mathbf{I}}-{\mathbf{Q}}_{(i)}({\mathbf{Q}}_{(i)}^{\top}{\mathbf{Q}}_{(i)})^{-1}{\mathbf{Q}}_{(i)}^{\top}$ and ${\mathbf{Q}}_{(i)}=({\mathbf{q}}_{1},\ldots,{\mathbf{q}}_{i-1})$. We control the norm of ${\mathbf{q}}_{i}$ by adding a constraint that the remainder has a norm of 1 if we project ${\mathbf{q}}_{i}$ on the space spanned by $\{{\mathbf{q}}_{1},\ldots,{\mathbf{q}}_{i-1}\}$, for $i=2,\ldots,r$. Note that the sparsity of ${\mathbf{q}}_{i}$ and ${\mathbf{a}}_{i}^{s}$ is the same. Compared to model (1), model (3) not only provides a clearer interpretation of factors but also alleviates the ambiguity issue.
@@
${\mathbf{A}}=\left(\begin{array}[]{ccc}1&1&1\\ 1&1&1\\ 0&1&1\\ 0&0&1\\ 1&2&2\end{array}\right),\quad{\mathbf{A}}^{s}=\left(\begin{array}[]{ccc}0&0&\sqrt{\frac{8}{3}}\\ 0&0&\sqrt{\frac{8}{3}}\\ 0&\sqrt{\frac{7}{2}}&0\\ \sqrt{3}&0&0\\ 0&\sqrt{\frac{7}{2}}&\sqrt{\frac{8}{3}}\\ \end{array}\right),\quad{\mathbf{Q}}=\left(\begin{array}[]{ccc}0&0&{\sqrt{2}}/{2}\\ 0&0&{\sqrt{2}}/{2}\\ 0&\sqrt{2}/2&0\\ 1&0&0\\ 0&\sqrt{2}/2&\sqrt{2}/2\\ \end{array}\right).$ (19)
@@
${\mathbf{V}}=\left(\begin{array}[]{ccc}0&0&1\\ 0&1&0\\ \frac{\sqrt{2}}{2}&\frac{\sqrt{2}}{2}&\frac{\sqrt{2}}{2}\end{array}\right).$
Hence, ${\mathbf{z}}_{t}$ in (3) satisfies ${\mathbf{z}}_{t}={\mathbf{V}}^{-1}{\mathbf{f}}_{t}$. ${\mathcal{M}}({\mathbf{A}})={\mathcal{M}}({\mathbf{A}}^{s})={\mathcal{M}}({\mathbf{Q}})$ but ${\mathbf{A}}^{s}$ and ${\mathbf{Q}}$ are more sparse, making the interpretation of the model much clearer. From the loading matrix ${\mathbf{Q}}$, we can see that only the 4-th time series loads on factor 1, the 3-rd and the 5-th time series loads on factor 2, and the 1-st, the 2-nd, and the 5-th time series load on factor 3.
Due to the latent nature of factors, various methods exist in the literature for separating factors from the noise process. Two assumptions are commonly used (Wang et al.,, 2019). The first one assumes that the factors have impacts on most of the time series, and thus the noise process can only accommodate weak cross-sectional dependence but is allowed to be weakly serially dependent; see Chamberlain and Rothschild, (1983); Bai and Ng, (2002); Bai, (2003); Stock and Watson, 2002b ; Bai and Ng, 2006a ; Hallin and Liska, (2007); Bai and Ng, (2008); Stock and Watson, (2010), among others. The second one assumes that factors capture all dynamics of the data and the noise process has no serial dependence, see Pan and Yao, (2008); Lam et al., (2011); Lam and Yao, (2012); Chang et al., (2015); Liu and Chen, (2016); Wang et al., (2019); Liu and Chen, (2022); Chen et al., (2022). In this paper, we adopt the second assumption, assuming that the error process is serially independent and uncorrelated with the factor process. Without loss of generality, we set the mean of the factor process to $\mathbf{0}$.

### v3

@@ Abstract
High-dimensional data analysis using traditional models suffers from overparameterization. Two types of techniques are commonly used to reduce the number of parameters – regularization and dimension reduction. In this project, we combine them by imposing a sparse factor structure and propose a regularized estimator to further reduce the number of parameters in factor models. One issue of factor analysis is that factors are hard to interpret, as both factors and the loading matrix are unobserved. To address this, we introduce a penalty term when estimating the loading matrix for a sparse estimate. As a result, each factor only drives a smaller subset of time series that exhibit the strongest correlation, improving the factor interpretability and predictive performance. Compared with existing work, the proposed method yields an estimator which enjoys the oracle property, and works in more general settings. The proposed method does not require the loading matrix to be orthogonal or factors to be independent. The theoretical properties of the proposed estimator are investigated. The simulation results are presented to confirm that our algorithm performs well. We apply our method to Hawaii tourism data.
Keywords: Regularization; dimension reduction; ADMM algorithm; factor models.
@@ 1 Introduction
High-dimensional time series data are widely observed in various disciplines, including finance, economics, business, and medical science. However, when analyzing high-dimensional data, traditional models, such as vector autoregressive models and regression models, suffer from overparameterization as noted in Lütkepohl (2005) and Chang et al. (2015). Consequently, traditional methods designed for these models are often not satisfactory in terms of prediction accuracy and model interpretability (Tibshirani, 1996). To address this problem, two types of methods are commonly used to reduce the number of parameters — regularization and dimension reduction (Liu and Zhang, 2022). Regularization methods resort to a sparsity assumption on the coefficient matrices and incorporate a penalty in the optimization function to obtain meaningful estimators; see examples Basu and Michailidis (2015) and Medeiros and Mendes (2016), while dimension reduction assumes that the high-dimensional data can be represented by a low-dimensional process. Among various dimension reduction tools, factor analysis is a popular choice and has been extensively investigated in the literature (Peña and Box, 1987; Li and Li, 1996; Bai and Ng, 2002; Stock and Watson, 2002a; Diebold and Li, 2006; Lam et al., 2011; Ando and Bai, 2017). In this paper, we combine both methods by imposing a sparse factor structure on the observed data and propose a regularized estimator for the loading matrix to further reduce the number of parameters. Our proposed method will not only enhance the model interpretation but also build a more parsimonious model to improve the predictive performance.
@@
Factor analysis has been an active research area in statistics and econometrics with a comprehensive theoretical framework (Bai and Ng, 2002; Bai, 2003; Lam et al., 2011; Lam and Yao, 2012) and numerous extensions, including factor models with structural breaks (Chen, 2015; Baltagi et al., 2017; Barigozzi et al., 2018; Baltagi et al., 2021), threshold mechanism (Massacci, 2017; Liu and Chen, 2020), regime switching (Liu and Chen, 2016), and constraints (Chen et al., 2020), and factor models for matrix data (Wang et al., 2019) and tensor data (Chen et al., 2022; Han et al., 2022, 2024a, 2024b). One issue of factor analysis is that factors are difficult to interpret as both factors and the loading matrix are latent. Ideally, observed variables can be used to proxy unobserved common factors. Bai and Ng (2006a) proposed statistical tests to evaluate the latent and observed factors. However, identifying suitable proxy variables is often challenging, especially when dealing with new data or data that are not well understood yet. Another solution is to interpret factors via loadings. In this paper, we incorporate a penalty function into the estimation of the loading matrix to shrink certain loadings to zero. This encourages each factor to load on a smaller subset of time series that exhibit the strongest correlation and thus enhances interpretability, thereby providing guidance on how to find meaningful proxy variables.
@@
Two widely used statistical methods aimed at improving the interpretability of factor models are particularly worth mentioning — varimax rotation and sparse principal component analysis (SPCA). It is well known that the loading matrix is not uniquely defined and can rotate in the loading space, which is obtained by principal component analysis (PCA) or other matrix decomposition methods. Varimax seeks the rotation that maximizes the variance of the squared elements in the loading matrix. Hence, loadings yielded by varimax are either very large in magnitude or close to 0. Compared to varimax, the method we propose has two key advantages. First, while varimax produces estimated loadings that are near-zero, our method ensures an estimate with zero elements, enhancing factor interpretability. Second, our method searches for the solution in a more general setting, yielding a more sparse loading matrix estimate. Unlike varimax, which keeps the loading space invariant, we allow the estimator to deviate slightly from the loading space obtained from PCA or its variants. SPCA is an extension of PCA (Zou et al., 2006) and requires the principal components to be orthogonal. In contrast to SPCA, our method relaxes this assumption, leading to a more sparse loading matrix and improving the interpretability of the factor model. Although PCA is commonly used as a tool to conduct factor analysis, the objective of factor analysis is distinct: it seeks to identify latent factors that explain the relationships among observed variables.
@@
In regression models, regularization methods, where some constraints on the parameters were added to solve an inference problem, are widely used to enhance model interpretability, reduce model variation, and thus improve prediction accuracy. LASSO proposed by Tibshirani (1996), uses an $L_{1}$ penalized likelihood for linear regression models to conduct variable selection. Subsequently, other sparsity-inducing penalties, such as smoothly clipped absolute deviation (SCAD) (Fan and Li, 2001) and minimax concave penalty (MCP) (Zhang, 2010), were also considered. Estimators obtained by SCAD and MCP both enjoy the oracle properties, meaning they perform as well as if the true underlying model were known in advance (Zou, 2006), which is theoretically appealing (Bertsimas et al., 2020). However, regularization methods have not been fully studied in the context of factor models. Kristensen (2017), Pelger and Xiong (2022), Uematsu and Yamagata (2022b), Uematsu and Yamagata (2022a), and Wu et al. (2026) studied the factor models with sparse loadings and most of them formulated the estimation as regularized optimization problems with an $L_{1}$-norm penalty term. Compared with the existing work, our proposed method has the following advantages. First, we use MCP penalty and the proposed estimator enjoys the oracle property, while their estimators do not. Our method with MCP penalty addresses two limitations of methods with $L_{1}$-norm penalty: estimation bias and model selection inconsistency. Secondly, our method works in more general settings, and yields a more sparse estimate, which builds a more parsimonious model to improve model interpretation and predictive performance. We relax the orthogonal assumption on the loading matrix and the assumption that factors are independent. As mentioned earlier, the loading matrix can rotate in the loading space. The existing work first selects a specific rotation, which makes columns of the loading matrix orthogonal and/or the factor covariance matrix an identity matrix, and then imposes the sparsity assumption on that particular loading matrix. However, these restrictions that factors are independent and the loading matrix is orthogonal are often too strong for empirical analysis. For example, the Fama-French three-factor models include an overall market factor and factors related to firm size (SMB) and book-to-market equity (HML). Empirical evidence shows that the correlation between the overall market factor and HML is significantly different from 0, which is -0.38 in Fama and French (1993), -0.622 in Durand et al. (2011), -0.30 in Fama and French (2015). Another example is the geographic location (i.e. Europe, Asia, Africa, etc) and the economic condition including the level of economic development (i.e. advanced vs emerging), which are considered as two important determinants for international business activities (Kose et al., 2003; Sanyal, 2005; Fernhaber et al., 2008). Since Europe has the most advanced economies and Africa has the most emerging economies, it is not appropriate to assume that the loadings of these two factors are orthogonal. In this paper, we relax these restrictions and explore the entire loading space to identify the most sparse loading matrix, which is more general and leads to a more sparse estimate. Numerical experiments and real data analysis in Section 5 and Section 6 show that our method outperforms the existing methods.
@@ 2 Model
We introduce some notations first. For a vector $\mathbf{z}$}, we use $z_{i}$ to denote its $i$-th element. For a $p_{1}\times p_{2}$ matrix $\mathbf{Z}$}, its ($i,j$)-th element is denoted by $z_{ij}$ and its $i$-th column is denoted by ${\mathbf{z}}_{i}$. Furthermore, we use ${\cal M}({\mathbf{Z}})$ to denote the space spanned by the columns of $\mathbf{Z}$}. Let $\|{\mathbf{Z}}\|_{F}$ be the Frobenius norm of $\mathbf{Z}$}, where $\|{\mathbf{Z}}\|_{F}=\sqrt{\sum_{i=1}^{p_{1}}\sum_{j=1}^{p_{2}}z_{ij}^{2}}$, $\|{\mathbf{Z}}\|_{2}$ be the LL-2 norm of $\mathbf{Z}$}, where $\|{\mathbf{Z}}\|_{2}=\left[\lambda_{\max}({\mathbf{Z}}^{\top}{\mathbf{Z}})\right]^{1/2}$ and $\lambda_{\max}(\cdot)$ is the maximum eigenvalue of a square matrix, and $\|{\mathbf{Z}}\|_{\min}$ is the nonzero minimum singular value of $\mathbf{Z}$}. We also define $_{1}, LL_{\infty}$, and max norm of a matrix as follows: $\|{\bf Z}\|_{1}=\max_{1\leq j\leq p_{2}}\sum_{i=1}^{p_{1}}|z_{ij}|$, $\|{\bf Z}\|_{\infty}=\max_{1\leq i\leq p_{1}}\sum_{j=1}^{p_{2}}|z_{ij}|$, and $\|{\bf Z}\|_{\max}=\max_{ij}|z_{ij}|$. We follow Cape et al. (2019) and define two-to-infinity norm as $\|{\bf z}\|_{\infty}=\max_{i}|z_{i}|$ for a vector $\mathbf{z}$} and as $\|{\bf Z}\|_{2\rightarrow\infty}=\sup_{\|{\mathbf{x}}\|_{2}=1}\|{\bf Z}{\bf x}\|_{\infty}$ for a matrix $\mathbf{Z}$}. For a scalar $z$, $|z|$ denotes its absolute value; for a set $\mathcal{Z}$, $|\mathcal{Z}|$ is the number of elements in the set; for a vector $\mathbf{z}$} or a matrix $\mathbf{Z}$}, $|{\mathbf{z}}|$ or $|{\mathbf{Z}}|$ is the number of its nonzero elements. For any $\{a_{n}\}$ and $\{b_{n}\}$, $a_{n}\asymp b_{n}$” means $\lim_{n\to\infty}a_{n}/b_{n}=c$, where $c$ is a positive constant, and $a_{n}\gtrsim b_{n}$” means $a_{n}^{-1}b_{n}=o(1)$.
@@
${\mathbf{x}}_{t}={\mathbf{A}}{\mathbf{f}}_{t}+\mbox{$\varepsilon$}_{t},$ (1)
where ${\mathbf{x}}_{t}$ is the $p$-dimensional time series, ${\mathbf{f}}_{t}=(f_{t1},f_{t2},\ldots,f_{tr})^{\top}$ is a set of unobserved (latent) factor time series with dimension $r$ that is much smaller than $p$, the matrix $\mathbf{A}$} is the loading matrix of the common factors, the term ${\mathbf{A}}{\mathbf{f}}_{t}$ can be viewed as the signal component of ${\mathbf{x}}_{t}$ and called the common component, and $\mbox{$\varepsilon$}_{t}$ is a noise process. The loading matrix $\mathbf{A}$} represents the impact of common factors ${\mathbf{f}}_{t}$ on the observed process ${\mathbf{x}}_{t}$.
@@
${\mathbf{x}}_{t}={\mathbf{A}}^{s}{\mathbf{f}}_{t}^{s}+\mbox{$\varepsilon$}_{t},$ (2)
@@
${\mathbf{x}}_{t}={\mathbf{Q}}{\mathbf{z}}_{t}+\mbox{$\varepsilon$}_{t},$ (3)
where ${\mathbf{q}}_{1}=\frac{{\mathbf{a}}_{1}^{s}}{\|{\mathbf{a}}_{1}^{s}\|_{2}}$ and ${\mathbf{q}}_{i}=\frac{{\mathbf{a}}_{i}^{s}}{\|{\mathbf{P}}_{i}{\mathbf{a}}_{i}^{s}\|_{2}}$, for $i=2,\ldots,r$, ${\mathbf{P}}_{i}={\mathbf{I}}-{\mathbf{Q}}_{(i)}({\mathbf{Q}}_{(i)}^{\top}{\mathbf{Q}}_{(i)})^{-1}{\mathbf{Q}}_{(i)}^{\top}$ and ${\mathbf{Q}}_{(i)}=({\mathbf{q}}_{1},\ldots,{\mathbf{q}}_{i-1})$. We control the norm of ${\mathbf{q}}_{i}$ by adding a constraint that the remainder has a norm of 1 if we project ${\mathbf{q}}_{i}$ on the space spanned by $\{{\mathbf{q}}_{1},\ldots,{\mathbf{q}}_{i-1}\}$, for $i=2,\ldots,r$. Note that the sparsity of ${\mathbf{q}}_{i}$ and ${\mathbf{a}}_{i}^{s}$ is the same. Compared to model (1), model (3) not only provides a clearer interpretation of factors but also alleviates the ambiguity issue.
@@
${\mathbf{A}}=\left(\begin{array}[]{ccc}1&1&1\\ 1&1&1\\ 0&1&1\\ 0&0&1\\ 1&2&2\end{array}\right),\quad{\mathbf{A}}^{s}=\left(\begin{array}[]{ccc}0&0&\sqrt{\frac{8}{3}}\\ 0&0&\sqrt{\frac{8}{3}}\\ 0&\sqrt{\frac{7}{2}}&0\\ \sqrt{3}&0&0\\ 0&\sqrt{\frac{7}{2}}&\sqrt{\frac{8}{3}}\\ \end{array}\right),\quad{\mathbf{Q}}=\left(\begin{array}[]{ccc}0&0&1\\ 0&0&1\\ 0&\sqrt{2}/2&0\\ 1&0&0\\ 0&\sqrt{2}/2&1\\ \end{array}\right).$ (19)
@@
${\mathbf{V}}=\left(\begin{array}[]{ccc}0&0&1\\ 0&\sqrt{2}&0\\ 1&1&1\end{array}\right).$
Hence, ${\mathbf{z}}_{t}$ in (3) satisfies ${\mathbf{z}}_{t}={\mathbf{V}}^{-1}{\mathbf{f}}_{t}$. ${\mathcal{M}}({\mathbf{A}})={\mathcal{M}}({\mathbf{A}}^{s})={\mathcal{M}}({\mathbf{Q}})$ but ${\mathbf{A}}^{s}$ and $\mathbf{Q}$} are more sparse, making the interpretation of the model much clearer. From the loading matrix $\mathbf{Q}$}, we can see that only the 4-th time series loads on factor 1, the 3-rd and the 5-th time series loads on factor 2, and the 1-st, the 2-nd, and the 5-th time series load on factor 3. Note that, the columns of $\mathbf{Q}$} are non-orthogonal in this example.
If the setting in Wu et al. (2026) is adopted, where the columns in the loading matrix are eigenvectors of $\mathbf{M}$} in (20) and are orthonormal, and for simplicity we further assume that $h_{0}=1$, $\{z_{t1}\}$, $\{z_{t2}\}$, and $\{z_{t3}\}$ are independent processes, and their auto-covariance matrix at lag 1 is identity, then the loading matrix will be written as ${\mathbf{v}}_{1},{\mathbf{v}}_{2},{\mathbf{v}}_{3}]$, where
${\mathbf{v}}_{1}=\frac{1}{\sqrt{30-10\sqrt{5}}}\left(\begin{array}[]{ccc}\sqrt{5}-1\\ \sqrt{5}-1\\ 3-\sqrt{5}\\ 0\\ 2\\ \end{array}\right),\quad{\mathbf{v}}_{2}=\left(\begin{array}[]{ccc}0\\ 0\\ 0\\ 1\\ 0\\ \end{array}\right),\quad{\mathbf{v}}_{3}=\frac{1}{\sqrt{30+10\sqrt{5}}}\left(\begin{array}[]{ccc}-\sqrt{5}-1\\ -\sqrt{5}-1\\ 3+\sqrt{5}\\ 0\\ 2\\ \end{array}\right).$
This loading matrix is less sparse than ours, $\mathbf{Q}$} in (19). From this example, we can see that our proposed method will reduce more parameters in factor models and build a more parsimonious factor model, which improves model interpretation and prediction accuracy.
Due to the latent nature of factors, various methods exist in the literature for separating factors from the noise process. Two assumptions are commonly used (Wang et al., 2019). The first one assumes that the factors have impacts on most of the time series, and thus the noise process can be weakly serially dependent and weakly cross-sectionally dependent; see Chamberlain and Rothschild (1983); Bai and Ng (2002); Bai (2003); Stock and Watson (2002b); Bai and Ng (2006b); Hallin and Liska (2007); Bai and Ng (2008); Stock and Watson (2010), among others. The second one assumes that factors capture all dynamics of the data and the noise process has no serial dependence but can accommodate strong cross-sectional dependence, see Pan and Yao (2008); Lam et al. (2011); Lam and Yao (2012); Chang et al. (2015); Liu and Chen (2016); Wang et al. (2019); Liu and Chen (2022); Chen et al. (2022). Among the existing studies on factor models with sparse loadings, Kristensen (2017), Pelger and Xiong (2022), Uematsu and Yamagata (2022b), and Uematsu and Yamagata (2022a) follow the first, while Wu et al. (2026) uses the second. In this paper, we adopt the second assumption, assuming no serial dependence in the noise process but allowing the strong cross-sectional dependence in the noise process. However, our method can be extended to cases where the first assumption holds.
Without loss of generality, we set the mean of the factor process to $\mathbf{0}$.